Wyle Company uses the weighted-average method in its process costing system. The following information was available for one of its processing departments:
• The beginning work in process inventory consisted of 16,000 units, which were 75% complete with respect to conversion costs.
• The department converted the equivalent of 59,000 units of production during the current month.
• A total of 50,000 units were completed and transferred out of that department during the current month.
• The ending work in process inventory consisted of 12,000 units, which were 65% complete with respect to conversion costs.
What is the number of equivalent units in ending work in process for the month for conversion costs? 10,000 or 7,800

Stay Corporation uses the weighted-average method in its process costing system. The company's ending work in process inventory consists of 8,000 units, 60% complete with respect to materials and 80% complete with respect to conversion costs. If the total cost in this inventory is $200,000 and if the cost for materials is $16 per equivalent unit for the period, the conversion cost per equivalent unit of the production for the period must be: 31.95 or 19.25 or None of the Above.

Parker Company uses a job-order costing system and applies manufacturing overhead to jobs using a predetermined overhead rate based on direct labor costs. Parker estimated 12000 machine hours. Last year manufacturing overhead and direct labor costs were estimated at $150,000 and 20,000, respectively, for the year. In July, Job #461 was completed. Materials costs on the job totaled $14,000 and labor costs totaled $1,500 at $5 per hour. At the end of the year, it was determined that the company incurred $ 124,000 in direct labor costs for the year and incurred $954,000 in actual manufacturing overhead costs.
Calculate the pre-determined overhead rate: 12.5 or 7.5 or None of the Above .
Answer #1

1) No Of equivalent unit under ending Work in process = 12000*65% = 7800

So answer is 7800 Units

2) Cost allocated to material = 8000*60%*16 = 76800

Conversion cost = 200000-76800 = 123200

Equivalent unit of conversion = 8000*80% = 6400

Cost per equivalent unit of conversion = 123200/6400 = 19.25

So answer is $19.25

3) Predetermine overhead rate = Estimated overhead/estimated labor cost = 150000/20000 = 7.50

So answer is 7.50